OpenStudy (2101vainilla):

PLEASE HELP!! WILL FAN AND MEDAL !!Sarah has two parties on the same day, one of them is from 7-10, while the other is from 7-12 The two venues are 30 minutes away from each other. At what time would Sarah have to leave the first party in order to make it to the second party and have been an equal amount of time at both?

OpenStudy (kodes):

Party 1- 3 Hours Party 2 - 5 Hours Distance between both parties- 30 minutes Total party time available = 12-7. Total party time = 5. In order to spend an equal amount of time at both, Total party time should be divided by 2. 5/2 = 2.5 Hours You can spend a maximum of 2.5 hours at each party. Time to travel from Party 1 to party 2 is 30 minutes. Half of 30 minutes is 15 minutes. Take 15 minutes from 2.5 hours. = 2 Hours and 15 Minutes. 7:00 + 2 Hours and 15 minutes. Therefore, Sarah would have to leave the first party at 9:15 in order to make it to the next party.

OpenStudy (2101vainilla):

Thank you so much!